OpenStudy (anonymous):

what is all groups of order up to 12?

OpenStudy (anonymous):

well there are a bunch of them

OpenStudy (anonymous):

prime order has to be cyclic, copies of \[\mathbb{Z_p}\]

OpenStudy (anonymous):

that takes care of 2,3,5, 7, 11

OpenStudy (anonymous):

two of order 4, namely Klien 4 aka \[\mathbb{Z_2}\times\mathbb{Z_2}\] and also \[\mathbb{Z_4}\]

OpenStudy (anonymous):

for order 6 you have \(S_3\) as well as \(\mathbb{Z_6}\)

OpenStudy (anonymous):

how about symmetric group and others group that have order 12

OpenStudy (anonymous):

ohh..thanks.. thats all?

OpenStudy (anonymous):

oh no there are plenty more. rather than me trying to remember them all, probably easiest to google

OpenStudy (anonymous):

order 8 there are 4 quaternions, \(\mathbb{Z_8}\) and \(\mathbb{Z_4}\times \mathbb{Z_2}\) and \(\mathbb{Z_2}\times \mathbb{Z_2}\times \mathbb{Z_2}\)

OpenStudy (anonymous):

let me find a link

OpenStudy (anonymous):

ok try here lists all i think http://en.wikipedia.org/wiki/List_of_small_groups

OpenStudy (anonymous):

thank you so much..^^